# The Loss of Control Playbook: Degrees, Dynamics, and Preparedness

### v1

@@ 2.1 Literature Reviewed
Lanham, Tamera, Anna Chen, Ansh Radhakrishnan, Benoit Steiner, Carson Denison. 2023. ”Measuring Faithfulness in Chain-of-Thought Reasoning.” https://arxiv.org/pdf/2307.13702.
@@
Baker, Bowen, Joost Huizinga, Leo Gao, Zehao Dou, Melody Y. Guan. 2025. ”Monitoring Reasoning Models for Misbehavior and the Risks of Promoting Obfuscation.” https://arxiv.org/abs/2503.11926.
@@
Hao, Shibo, Sainbayar Sukhbaatar, DiJia Su, Xian Li, Zhiting Hu. 2025. ”Training Large Language Models to Reason in a Continuous Latent Space.” https://arxiv.org/pdf/2412.06769.
@@
van der Weij, Teun, Felix Hofstätter, Ollie Jaffe, Samuel F. Brown, Francis Rhys Ward. 2025. ”AI Sandbagging: Language Models can Strategically Underperform on Evaluations.” https://arxiv.org/abs/2406.07358.
@@
Greenblatt, Ryan, Carson Denison, Benjamin Wright, Fabien Roger, Monte MacDiarmid. 2024. ”Alignment faking in large language models.” https://arxiv.org/abs/2412.14093.
@@
Finnveden, Lukas, C. Jess Riedel, Carl Shulman. 2022. ”AGI and Lock-in.” https://www.forethought.org/research/agi-and-lock-in.
Anthropic. 2025. ”Responsible Scaling Policy.” https://www-cdn.anthropic.com/872c653b2d0501d6ab44cf87f43e1dc4853e4d37.pdf.
@@
Shlegeris, Buck, Fabien Roger, Lawrence Chan, Euan McLean. 2024. ”Language models are better than humans at next-token prediction.” https://arxiv.org/pdf/2212.11281.
@@
Nevo, Sella, Dan Lahav, Ajay Karpur, Yogev Bar-On, Henry Bradley. 2024. ”Securing AI Model Weights: Preventing Theft and Misuse of Frontier Models.” https://www.rand.org/pubs/research˙briefs/RBA2849-1.html.

### v2

@@ 2.1 Literature Reviewed
Lanham, Tamera, Anna Chen, Ansh Radhakrishnan, Benoit Steiner, Carson Denison. 2023. “Measuring Faithfulness in Chain-of-Thought Reasoning.” https://arxiv.org/pdf/2307.13702.
@@
Baker, Bowen, Joost Huizinga, Leo Gao, Zehao Dou, Melody Y. Guan. 2025. “Monitoring Reasoning Models for Misbehavior and the Risks of Promoting Obfuscation.” https://arxiv.org/abs/2503.11926.
@@
Hao, Shibo, Sainbayar Sukhbaatar, DiJia Su, Xian Li, Zhiting Hu. 2025. “Training Large Language Models to Reason in a Continuous Latent Space.” https://arxiv.org/pdf/2412.06769.
@@
van der Weij, Teun, Felix Hofstätter, Ollie Jaffe, Samuel F. Brown, Francis Rhys Ward. 2025. “AI Sandbagging: Language Models can Strategically Underperform on Evaluations.” https://arxiv.org/abs/2406.07358.
@@
Greenblatt, Ryan, Carson Denison, Benjamin Wright, Fabien Roger, Monte MacDiarmid. 2024. “Alignment faking in large language models.” https://arxiv.org/abs/2412.14093.
@@
Finnveden, Lukas, C. Jess Riedel, Carl Shulman. 2022. “AGI and Lock-in.” https://www.forethought.org/research/agi-and-lock-in.
Anthropic. 2025. “Responsible Scaling Policy.” https://www-cdn.anthropic.com/872c653b2d0501d6ab44cf87f43e1dc4853e4d37.pdf.
@@
Shlegeris, Buck, Fabien Roger, Lawrence Chan, Euan McLean. 2024. “Language models are better than humans at next-token prediction.” https://arxiv.org/pdf/2212.11281.
@@
Nevo, Sella, Dan Lahav, Ajay Karpur, Yogev Bar-On, Henry Bradley. 2024. “Securing AI Model Weights: Preventing Theft and Misuse of Frontier Models.” https://www.rand.org/pubs/research˙briefs/RBA2849-1.html.